# Mitigating Low-Frequency Bias: Feature Recalibration and Frequency Attention Regularization for Adversarial Robustness

### v1

@@ IV-B1 Comparison on CIFAR-10 and CIFAR-100
(1) HFDR improves the robustness of different adversarial training methods. After integrating our proposed HFDR module into the three baselines, the HFDR-AT, HFDR-MART, and HFDR-AWP can consistently outperform their corresponding baseline models across all attack scenarios, showcasing superior performance. For example, by incorporating HFDR into the PGD-AT method, we observe performance improvements of 2.60% in PGD-10, 1.22% in CW, and 2.07% in AA on CIFAR-10. Additionally, there is an increase in clean accuracy by 0.97% on CIFAR-100. Moreover, integrating the HFDR module into the MART and AWP methods also results in significant improvements in robust performance. The performance enhancement can be attributed to extracting valuable hidden semantics from high-frequency features that contain concentrated disturbances.
(2) Our model outperforms the baseline under attack. Our proposed model, HFDR-AWP, exhibits superior performance compared to all baseline methods, showcasing exceptional robustness. Specifically, HFDR-AWP outperforms the previous best baseline AWP by 1.77% of PGD-10, 1.51% of PGD-50, and 2.34% of AA on CIFAR-100. Besides, our method significantly outperforms the state-of-the-art frequency-based method, FPCM by a large margin. Unlike FPCM, which inserts multiple blocks into the network, we only inserted a single block, reducing the additional computation cost.
@@
IV-B2 Comparison on Tiny ImageNet
To evaluate the credibility and generalizability of the performance improvements, the corresponding performance results on the Tiny ImageNet [48] are presented in Table II. This dataset features higher resolution and a wider range of categories compared to CIFAR-10 and CIFAR-100. Specifically, the evaluation results on the Tiny ImageNet dataset reveal that HFDR-AT achieved performance enhancements of 1.86% and 2.11% against C&W and AA attacks, respectively, compared to PGD-AT. Furthermore, HFDR-AWP exhibited improvements of 3.47% and 2.30% when compared to AWP. These findings suggest that our approach seamlessly integrates with adversarial training frameworks, maintaining robust performance when confronted with complicate datasets.
| Method | Clean | PGD-10 | PGD-20 | PGD-50 | C&W | AA |
| PGD-AT [31] | 44.15 | 21.45 | 21.08 | 20.91 | 18.74 | 16.13 |
| MART [33] | 45.13 | 23.41 | 22.37 | 22.17 | 18.91 | 16.32 |
| AWP [34] | 44.86 | 22.03 | 21.72 | 21.50 | 19.08 | 16.97 |
| HFDR-AT | 45.48 | 24.68 | 23.80 | 23.14 | 20.60 | 18.24 |
| HFDR-MART | 46.08 | 25.71 | 25.05 | 24.47 | 21.19 | 18.97 |
| HFDR-AWP | 45.67 | 26.69 | 25.28 | 24.61 | 22.55 | 19.27 |
@@ IV-F Parameter Sensitive Analysis
In this part, we performed an analysis of the HFDR module hyper-parameters. In Figure 6, a visualization of parameter sensitivity for hyperparameters and in the HFDR method is demonstrated. Notably, assigning a value of zero to results in the absence of the frequency attention regularization term (refer Equation 10). Evidence suggests that these two hyper-parameters influence the trade-off between model clean and robust accuracy performance. The following offers a detailed analysis:
@@ IV-G Analysis of Mitigation of Low-Frequency Bias
In this part, we delve into exploring the mitigating effects of our proposed module HFDR on low-frequency biases during adversarial training. In Figure 5, we showcased the efficacy of our HFDR-AT method in extracting diverse frequency domain features compared to PGD-AT. The left figure illustrates that the growth trend in model performance with the incorporation of low to high-frequency features, whereas the right figure depicts variations in model performance with the incorporation of low to high-frequency components. These findings suggest that our model exhibits an enhanced proficiency in extracting high-frequency features. Specifically, our model can recalibrate the perturbed high-frequency features to alleviate the low-frequency bias caused by traditional adversarial training.
@@
IV-H Computational Efficiency
Table VIII depicts the comparison of our approach and the original model in terms of training efficiency. This comparison includes the analysis of the model parameters and the average time required for one training cycle. The results highlight that our model enhances the robustness of the model without a significant increase in model parameters or the cost of AT. For instance, on the CIFAR-10, the regular WRN34-10 takes 1,050 seconds for one training cycle, whereas the model with HFDR only takes 1,068 seconds. This demonstrates the clear advantages of our model compared to other methods with complex network architectures.
@@ V Conclusion
In this paper, we propose a module called High-Frequency Disentanglement and Recalibration (HFDR) to tackle the issue of low-frequency bias arising from adversarial training, which comprises three key components: feature disentanglement and recalibration with frequency domain attention regularization. We achieve this by leveraging frequency domain attention maps to separate feature maps into low-frequency and high-frequency components and then recalibrating the high-frequency features. This recalibration enhances the model’s robustness by incorporating valuable semantic clues for accurate prediction. Additionally, we utilize frequency domain attention regularization to ensure a balanced extraction of features with different frequencies during the derivation process, thereby mitigating the low-frequency bias associated with traditional adversarial training. Extensive experiments verified the effectiveness of our method.

### v2

@@ IV-B1 Comparison on CIFAR-10 and CIFAR-100
(1) HFDR improves the robustness of different adversarial training methods. After integrating our proposed HFDR module into the three baselines, the HFDR-AT, HFDR-MART, and HFDR-AT-AWP can consistently outperform their corresponding baseline models across all attack scenarios, showcasing superior performance. For example, by incorporating HFDR into the PGD-AT method, we observe performance improvements of 2.60% in PGD-10, 1.22% in C&W, and 2.07% in AA on CIFAR-10. Additionally, there is an increase in clean accuracy by 0.97% on CIFAR-100. Moreover, integrating the HFDR module into the MART and AWP methods also results in significant improvements in robust performance. The performance enhancement can be attributed to extracting valuable hidden semantics from high-frequency features that contain concentrated disturbances.
| Method | Clean | PGD-10 | PGD-20 | PGD-50 | C&W | AA |
| PGD-AT [33] | 83.80 | 58.11 | 57.56 | 57.08 | 55.35 | 54.69 |
| MART [35] | 81.17 | 59.71 | 59.23 | 59.03 | 57.17 | 56.18 |
| AWP [36] | 83.25 | 61.32 | 60.92 | 60.74 | 59.39 | 57.05 |
| HFDR-AT | 85.39 | 60.59 | 60.24 | 59.46 | 58.10 | 57.55 |
| HFDR-MART | 86.22 | 62.50 | 62.18 | 61.84 | 60.71 | 59.56 |
| HFDR-AT-AWP | 84.95 | 62.94 | 62.77 | 62.41 | 61.38 | 60.67 |
(2) Our model outperforms the baseline under attack. Our proposed model, HFDR-AT-AWP, exhibits superior performance compared to all baseline methods, showcasing exceptional robustness. Specifically, HFDR-AT-AWP outperforms the previous best baseline AWP by 1.77% of PGD-10, 1.51% of PGD-50, and 2.34% of AA on CIFAR-100. Besides, our method significantly outperforms the state-of-the-art frequency-based method, FPCM by a large margin. Unlike FPCM, which inserts multiple blocks into the network, we only inserted a single block, reducing the additional computation cost.
@@
IV-B2 Comparison on Tiny ImageNet
To evaluate the credibility and generalizability of the performance improvements, the corresponding performance results on the Imagenette [56] are presented in Table II. This dataset features higher resolution compared to CIFAR-10 and CIFAR-100. Specifically, the evaluation results on the Imagenette dataset reveal that HFDR-AT achieved performance enhancements of 2.75 and 2.86% against C&W and AA attacks, respectively, compared to PGD-AT. Furthermore, HFDR-AT-AWP exhibited improvements of 1.99% and 3.61% when compared to AWP. These findings suggest that our approach seamlessly integrates with adversarial training frameworks, maintaining robust performance when confronted with complicate datasets.
@@ IV-F Parameter Sensitive Analysis
In this part, we performed an analysis of the HFDR module hyper-parameters. In Figure 7, a visualization of parameter sensitivity for hyperparameters and in the HFDR method is demonstrated. Additionally, we also analyzed the impact of and (HFDR-MART) in Figure 8. Notably, assigning a value of zero to results in the absence of the frequency attention regularization term (refer Equation 11). Evidence suggests that these two hyper-parameters influence the trade-off between model clean and robust accuracy performance. The following offers a detailed analysis:
@@
(3) The parameter controls the trade-off between clean accuracy and robust accuracy. In the left figure, we fixed and analyzed how varying affects both clean accuracy and robustness. The right figure assesses the impact of varying while keeping . For these two figures, we find that increasing or will generally decrease clean accuracy while increasing robustness. To have a balance, we choose (following [18, 49, 6]) and in this study to have a balance between clean accuracy and robust accuracy.
| Dataset | Method | Clean | PGD-10 | PGD-20 | PGD-50 | C&W | AA |
| CIFAR-10 | HFDR-AT | 85.89 ( 0.34) | 58.67 ( 0.62) | 57.69 ( 0.51) | 56.54 ( 0.47) | 55.13 ( 0.40) | 53.76 ( 0.38) |
| HFDR-MART | 83.43 ( 0.41) | 59.73 ( 0.71) | 58.96 ( 0.63) | 58.73 ( 0.56) | 54.56 ( 0.49) | 53.92 ( 0.42) | |
| HFDR-AT-AWP | 85.19 ( 0.30) | 60.46 ( 0.55) | 59.28 ( 0.44) | 58.76 ( 0.33) | 56.62 ( 0.28) | 54.64 ( 0.22) |
@@ IV-G Analysis of Mitigation of Low-Frequency Bias
In this part, we delve into exploring the mitigating effects of our proposed module HFDR on low-frequency biases during adversarial training. In Figure 5, we showcased the efficacy of our HFDR-AT method in extracting diverse frequency domain features compared to PGD-AT. The left figure illustrates that the growth trend in model performance with the incorporation of low to high-frequency features, whereas the right figure depicts variations in model performance with the incorporation of low to high-frequency components. Additionally, as visualized in Figure 9, our method significantly enhances the extraction of high-frequency features while maintaining the ability to capture low-frequency features. These results indicate that our model recalibrates vulnerable high-frequency features, mitigating the low-frequency bias inherent in traditional adversarial training.
IV-H Evaluation of Experimental Stability
To evaluate the stability of our experimental results, we performed five independent runs of each experiment and reported the mean and standard deviation of the outcomes. As shown in Table VIII, the results exhibit minimal fluctuations across multiple runs, confirming the robustness and consistency of the HFDR method under varying conditions.
IV-I Computational Efficiency
Table IX depicts the comparison of our approach and the original model in terms of training efficiency. This comparison includes the analysis of the model parameters and the average time required for one training cycle. The results highlight that our model enhances the robustness of the model without a significant increase in model parameters or the cost of AT. For instance, on the CIFAR-10, the regular WRN34-10 takes 1,050 seconds for one training cycle, whereas the model with HFDR only takes 1,068 seconds. This demonstrates the clear advantages of our model compared to other methods with complex network architectures.
@@ V Conclusion
In this paper, we presented a comprehensive analysis of frequency-domain characteristics in adversarially trained networks and introduced HFDR (High-Frequency Disentanglement and Recalibration), a novel approach to enhance model robustness. Our investigation revealed a critical low-frequency bias in conventional adversarial training methods, where models predominantly rely on low-frequency components while overlooking valuable semantic information contained in high-frequency features. To address this limitation, we developed the HFDR module, which employs high-pass filters and frequency attention mechanisms to effectively disentangle and recalibrate high-frequency features. Through the integration of frequency domain attention regularization, our approach successfully harmonizes feature extraction across different frequency domains, mitigating the inherent low-frequency bias of adversarial training. Extensive experiments verified the effectiveness of our method.